OpenStudy (anonymous):

how many grams of hydrogen in 4.50 mol H2SO4?

OpenStudy (anonymous):

H=1 gr/mol s=32 gr/mol O=16 gr/mol and with using them we want to calculate molar mass of H2SO4 H2SO4=2+32+(4*16)=98 gr/mol then you can use this formula to find the mass of H2SO4 n=m/M n=4.5 mol M=98 gr/mol m=? gr

OpenStudy (anonymous):

@gingerrx

OpenStudy (anonymous):

\[n=\frac{ m }{ M }\]\[4.5=\frac{ m }{ 98 }\]\[m=4.5\times 98=441\]

OpenStudy (anonymous):

given= unknown divided by molar mass?

OpenStudy (anonymous):

? means we should calculate m with n=4.5 and M=98 gr/mol but he haven't find the mass of hydrogen.we find it with using this formula molar mass of hydrogen/molar mass of H2SO4=mass of hydrogen/ mass of H2SO4

OpenStudy (anonymous):

i was clarifying, thank you!!

OpenStudy (anonymous):

this is the final step \[\frac{ 2 }{ 98 }=\frac{m _{h}}{ 441 }\]\[m _{h}=\frac{ 441\times 2 }{ 98 }=9 gr\]

OpenStudy (anonymous):

woo awesome, thanks i finally get it